SkyChefs, Inc., prepares in-flight meals for a number of major airlines. One of the company’s products is grilled salmon in dill sauce with baby new potatoes and spring vegetables. During the most recent week, the company prepared 4,800 of these meals using 2,350 direct labor-hours. The company paid its direct labor workers a total of $23,500 for this work, or $10.00 per hour. According to the standard cost card for this meal, it should require 0.50 direct labor-hours at a cost of $9.40 per hour. Required: 1. What is the standard labor-hours allowed (SH) to prepare 4,800 meals? 2. What is the standard labor cost allowed (SH × SR) to prepare 4,800 meals? 3. What is the labor spending variance? 4. What is the labor rate variance and the labor efficiency variance? (For requirements 3 and 4, indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values. Do no round intermediate calculations.)
Answer #1

Standard hours allowed = 4800*0.50

= 2400

Standard cost allowed = 2400*9.40

= 22,560

Labor spending variance = 22,560 - 23,500

= 940 U

Labor Rate Variance = (SR-AR) *AH

= (9.40 - 10)*2350

= 1410 U

Labor efficiency variance = (SH-AH) *SR

= (2400-2350)*9.40

= 470 F